OpenStudy (anonymous):

Write and solve an inequality. An airline requires carry-on luggage to weigh at most 40 pounds. Your suitcase currently weighs 10 pounds. How many pounds p are available for you to fill your suitcase with other items?

OpenStudy (ajanijones):

You said there are 30 pounds left over and the question tells you p are the available pounds left to fill your suitcase. So p = 30. You know this. 40 - 30 = 10 which is not less than 10, so we can discount the one you thought it was. If the inequality said 40 - 30 <= 10 , fine - we couldn't, but it doesn't. Now, let's look the bottom inequality. 10 + p <= 40 ... Well this is true 10 + 30 = 40 which is less than or equal to 40. p <= 30 is also true, since p is equal to 30.

OpenStudy (ajanijones):

Note: the third inequality is actually true too, and you should find this if you have understood what I have said

OpenStudy (ajanijones):

It is both.